EEMC HV adjustments for "outliers"


EEMC HV adjustments for "outliers"

Using the sums method (similar to Run 9 analysis) Scott Identified some outliers who's gains were either too high or too low compared to other towers in the same eta ring.  The list of towers is below in two different groups 1) known bad channels and 2) channels to adjust HV.

1) Known bad channels:

06TA03, 02TC06, 07TC05 -> all reported to still be bad and masked at L0

04TB05 -> spectra shows this channel is dead at startup

2) Channels to adjust HV

    gain too high:  11TB08, 03TA06, 08TC03, 07TC07, 02TE02, 01TA04

    gain too low:  12TD01, 10TA09



Minbias runs (12030069-73) were used to determine the slope of each channel using "Scott's method".  Below is a summary of the channels slope and the median slope for towers in that eta bin as well as the current HV (HVset_ix) for each tower, all of which is used to calculate the new HV (HVset_xi) to be used to match the other towers in that eta bin.

Table 1:

The equation used to determine the new HV values is HV_1 = HV_2 * (slope_2 / slope_1) ^ (1/kappa), where the value of kappa is taken to be 8.8 from previous HV adjustments.    This is equation is different than the similar equation used by the barrel becuase in the endcap gain ~ 1/slope and in the barrel gain ~ slope.


1) 11TB08 appears to be unstable (HV status is 5 or 7 instead of 4="good") running at 517 V, so it was decided to set this tower to its original value of 699.9 V where it runs stably.  Thus, this tower will continue to be masked out of L0 and L2 trigger and will need to be calibrated carefully to be used in offline analysis.

2) 12TD01 is a bit of a special case as it was "hot" in Run 8 and had its voltage lowered (946.2 -> 830.8) before Run 9, but then the gain was way too low.  So it was decided to put its voltage at 880.0 to try and increase the gain without getting hot.


Runs taken to test new HV

Once new HV values were determined 2 emc-check runs were taken to check the new files.  12037043 was taken with HVset_ix and 12037046 was taken with HVset_xi.  Below is a summary of the slopes for the 2 runs for the channels of interest.  Unfortunately. some channels didn't get a new HV loaded because of communications problems with HVsys branch "C" (these are shown in blue). 

Table 2:

For the towers where the new HV was loaded correctly (red) the slopes now match much better to the median slope in it's etabin, so the new HV values look reasonable and will be used for the remainder of Run 11.  HVset_xi is used for all runs after R12038072.

Note:  10TA09 appeared to be hot after its HV was raised to 827.0 so it was set back to its HVset_ix voltage value.  There was probably beam background when the original slopes were measured causing this to incorrectly be labeled an outlier.

5P1 adjustments

During the same test runs (12037043 and 12037046) a test was done increasing tube #2 of 5P1 from 750 -> 840 V, as the spectra for channels 176-191 was way down in some early runs (eg. 12034091).  This increased the gain by ~2.6, but comparisons of the slopes for channels from tube #2 to other preshower channels in 5P1 showed that the gain should be increased by another factor of ~2 (see txt file with slopes for tubes of interest and median slope of other channels in run 12037046). 

So the final HV used for 5P1 will be 913 V, which was determined with a similar formula as for the towers, but with kappa_mapmpt=8.3.